# Enhanced Curvature Perturbation and Primordial Black Hole Formation in Two-stage Inflation with a break

### v1

@@ 1 Introduction
PBHs contain rich phenomena of physics. First, since PBHs interact with other matter only via gravity, and they are born static, they are a natural candidate for cold dark matter (CDM) [16]. In particular, for asteroid mass (ranging from $10^{-16}M_{\odot}$ to $10^{-12}M_{\odot}$) PBHs, they may constitute all the CDM [17, 18, 19, 20]. Second, PBHs also provide a natural explanation for the high masses and low spins of binary black hole (BBH)  [21, 22, 23, 24, 25, 26, 27, 28, 29], inferred from the past three runs of GW observations by LIGO-Virgo and LIGO-Virgo-KAGRA collaborations [30, 31, 32]. The recently discovered signal of a GW background by several pulsar timing array (PTA) collaborations at $10^{-9}\sim 10^{-8}$ Herz band [33, 34, 35, 36, 37, 38, 39, 40, 41] indicates the existence of sub-solar mass PBHs, which may be detected by microlensing observations [42, 43, 44, 45, 46, 47, 48, 49, 17].
There are a number of proposed mechanisms to enhance the primordial curvature perturbations during inflation, for instance, single-field inflation with a non-slow-roll stage  [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70], multi-field inflation [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90], modify gravity [91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120], curvaton scenario [121, 122, 123, 124, 125, 126], parametric resonance [127, 128, 129, 130], etc. The property that any successful model must have is its consistency with the CMB and LSS constraints on large scales. In other words, it is essential to decouple the physics on small scales from that on large scales. This can be naturally realized in the scenario where inflation consists of two or more distinct periods intervened by a transition stage that may or may not break inflation. We call it the two-stage inflation. It was proposed by various people for various purposes [131, 132, 133, 134, 135].
In this paper, we investigate the enhancement of primordial curvature perturbations and the PBH formation in the two-stage inflationary scenario. As a concrete model, we adopt Starobinsky’s $R^{2}$ gravity theory [136] supplemented with a non-minimally coupled scalar field, where the first stage is dominated by the scalaron of the $R^{2}$ theory while the second stage by the non-minimally coupled scalar field. The main reason for this choice is that it is easy to satisfy the observational constraint on the CMB scales while giving us enough degrees of freedom to design the second stage.
@@ 2 The model
$S_{J}=\int d^{4}x\sqrt{-g}\ \bigg{[}\frac{M_{\mathrm{pl}}^{2}}{2}f(R)-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\partial_{\nu}\chi-V(\chi)\bigg{]},$ (2.1)
@@
the action (2.1) can be expressed in the Einstein frame as [137, 138, 139, 140]
@@
Let us now comment on the role of $\chi_{0}$ in (2.2). As explained above, or as shown in Fig. 1, $\chi$ is trapped at the local minimum $\chi\approx\chi_{0}$ during St-1. If $\chi_{0}=0$, as soon as St-2 starts at which $\phi$ settles down to the local minimum $\phi=\phi_{c}$, the point $\chi=0$ becomes a local maximum, implying that classically $\chi$ would be stuck at $\chi=0$, and it would move away from $\chi=0$ only by the quantum fluctuations. This leads to unacceptably large curvature perturbations. A non-vanishing $\chi_{0}$ avoids this catastrophe.
Before closing this section, let us briefly comment on the stability of the trajectory during St-1. During most of this stage, we have $F\gg 1$ so that the effective mass of $\chi$ is positive and $\chi$ acts as a damped harmonic oscillator (See Fig. 1). In our model, we find that the coupling parameter $\xi$ must satisfy $\xi>3/16$ to ensure that $\chi$ undergoes an under-damped oscillators during St-1 so that it settles down to the stable local minimum $\chi_{0}$ in a few e-folds. This secures that St-1 can be effectively treated as single-field slow-roll inflation.
@@ 4.3 Superhorizon evolution and $\delta N$
$\mathcal{P}_{\mathcal{R}}(k)\approx\frac{M^{2}}{16\pi^{2}M_{\rm pl}^{2}}\times\left\{\begin{aligned} &\quad\left[\frac{2}{3}\left(\ln\frac{k_{1}}{k}+\frac{3}{4}F_{\star 1}\right)^{2}+g_{1}^{2}h^{2}\left(\frac{M_{\rm pl}}{\chi_{0}}\right)^{2}\frac{F_{\star 2}H_{\rm exit}}{H_{\star 2}}\left(\frac{k}{k_{1}}\right)^{3}\right]\\ &\qquad\times\left[1-\left(\frac{4}{3}\ln\frac{k_{1}}{k}+F_{\star 1}\right)^{-1}\right]^{2}\quad\mathrm{for}\ k<k_{1}\,,\\ &\quad\frac{g_{2}^{2}h^{2}}{\mu^{2}}\left(\frac{M_{\rm pl}}{\chi_{0}}\right)^{2}\left(\frac{k}{k_{2}}\right)^{-\beta}\quad\mathrm{for}\ k>k_{1}\,,\end{aligned}\right.$ (4.31)
where we have used $H_{\rm exit}^{2}=M^{2}(1-F_{\rm exit}^{-1})^{2}/4$ during St-1, $H^{2}=V_{0}/(3M_{\mathrm{pl}})^{2}$ during St-3, and introduced
@@ 5.1 Primordial black hole formation
Figure 12: Three numerical result of power spectrum $\mathcal{P}_{\mathcal{R}}$ and $f_{\mathrm{PBH}}$. Left panel: We set $\mathcal{P}_{\mathcal{R}}=2.3\times 10^{-9}$ at pivot scale $k=0.002\mathrm{Mpc^{-1}}$ according to [1]. From right to left, the green, blue, and pink curves are numerically calculated with parameter cases 1,2,3 appearing in Table. 1. We tune the parameters and obtain case 1 for PBH as a dominant candidate of cold dark matter, case 2 for the production of sub-solar PBH, and case 3 for explanation of the recently released PTAs observation intentionally. Right panel: From right to left, the green, blue, and pink curves are numerically calculated with parameter case 1,2,3 in Table.1. The corresponding power spectrum is shown by the same color in the left panel. The $f_{\rm PBH}$ constraint plot and codes are available at PBH bounds, we include the observational constraints from PBH evaporation [144] (evap), HSC [145]/EROS [146]/OGLE [17]/MACHO [147]/Kepler [148] microlensing (HSC/EROS/OGLE/M/K) observations and millilensing observation of compact radio sources [149] (Radio), dynamical heating of ultra-faint dwarf galaxies [150] (UFdwarfs), the accretion constraints by CMB anisotropies measured by Planck [151] (CMB), and background constraints by LIGO [152, 153, 154] (LIGO/LIGO-SGWB).
@@
$\chi_{0}/M$ $0.275\sqrt{1/(2\pi^{2})}$ $0.172\sqrt{1/(2\pi^{2})}$ $0.141\sqrt{1/(2\pi^{2})}$
@@ 6 Conclusion and Discussion
In terms of the $\delta N$ formalism, this enhancement may be understood as a consequence of the sharp turn of the inflationary trajectory in the field space, an almost orthogonal turn from the $\phi$-direction to the $\chi$-direction. This converts the original isocurvature perturbation due to $\delta\chi$ at St-1 to an adiabatic perturbation at St-3. Since $\delta\chi$ decays like $\sim a^{-3}$ during St-1, the isocurvature effect is negligible at early stages. Conversely, it means that the isocurvature effect grows like $a^{3}$ toward the end of St-1. This leads to the growth of the curvature perturbation spectrum proportional to $k^{3}$. This is our new finding. This growth rate is universal as long as the isocurvature field ($\chi$ in our case) is heavy enough at the first inflationary stage and dominates the second inflationary stage as a slow-rolling inflaton.
@@
In this paper, we assumed that the curvature perturbation is Gaussian. However, it is known that non-Gaussianities may have a significant influence on the PBH abundance [163, 164, 165] as well as on the PBH clustering [166, 167, 168], which may also leave characteristic features on the corresponding induced GWs [169, 65, 170, 171, 172, 173]. In our model, we found that the non-Gaussianities are of local type, and those generated around the peak of the power spectrum are very small, as expected in the case of multi-field inflationary scenarios with non-extreme transitions [155, 156].

### v2

@@ 1 Introduction
PBHs contain rich phenomena of physics. First, since PBHs interact with other matter only via gravity, and they are born static, they are a natural candidate for cold dark matter (CDM) [16]. In particular, for asteroid mass (ranging from $10^{-16}M_{\odot}$ to $10^{-12}M_{\odot}$) PBHs, they may constitute all the CDM [17, 18, 19, 20]. Second, PBHs also provide a natural explanation for the high masses and low spins of binary black hole (BBH)  [21, 22, 23, 24, 25, 26, 27, 28, 29], inferred from the past three runs of GW observations by LIGO-Virgo and LIGO-Virgo-KAGRA collaborations [30, 31, 32]. The recently discovered signal of a GW background by several pulsar timing array (PTA) collaborations at $10^{-9}\sim 10^{-8}$ Herz band  [33, 34, 35, 36] has drawn a widespread attention [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], it indicates the existence of sub-solar mass PBHs, which may be detected by microlensing observations [48, 49, 50, 51, 52, 53, 54, 55, 17].
There are a number of proposed mechanisms to enhance the primordial curvature perturbations during inflation, for instance, single-field inflation with a non-slow-roll stage  [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80], multi-field inflation [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102], modify gravity [103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132], curvaton scenario [133, 134, 135, 136, 137, 138], parametric resonance [139, 140, 141, 142], etc. The property that any successful model must have is its consistency with the CMB and LSS constraints on large scales. In other words, it is essential to decouple the physics on small scales from that on large scales. This can be naturally realized in the scenario where inflation consists of two or more distinct periods intervened by a transition stage that may or may not break inflation. We call it the two-stage inflation. It was proposed by various people for various purposes [143, 144, 145, 146, 147, 148].
In this paper, we investigate the enhancement of primordial curvature perturbations and the PBH formation in the two-stage inflationary scenario. As a concrete model, we adopt Starobinsky’s $R^{2}$ gravity theory first proposed in [149] supplemented with a non-minimally coupled scalar field, where the first stage is dominated by the scalaron of the $R^{2}$ theory while the second stage by the non-minimally coupled scalar field. The main reason for this choice is that it is easy to satisfy the observational constraint on the CMB scales while giving us enough degrees of freedom to design the second stage.
@@ 2 The model
$S_{J}=\int d^{4}x\sqrt{-g}\ \bigg{[}\frac{M_{\mathrm{pl}}^{2}}{2}f(R)-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi-V(\chi)\bigg{]},$ (2.1)
@@
the action (2.1) can be expressed in the Einstein frame as [150, 151, 152, 153, 154, 155]
@@
Let us now comment on the role of $\chi_{0}$ in (2.2). As explained above, or as shown in Fig. 1, $\chi$ is trapped at the local minimum $\chi\approx\chi_{0}$ during St-1. If $\chi_{0}=0$, as soon as St-2 starts at which $\phi$ settles down to the local minimum $\phi=\phi_{c}$, the point $\chi=0$ becomes a local maximum, implying that classically $\chi$ would be stuck $\chi=0$, and it would move away from $\chi=0$ only by the quantum fluctuations. This leads to unacceptably large curvature perturbations. A non-vanishing $\chi_{0}$ avoids this catastrophe. The same strategy is also taken in [156] to avoid the overproduction of topological defects which often happens in hybrid inflation scenarios such as [82, 86]. Before closing this section, let us briefly comment on the stability of the trajectory during St-1. During most of this stage, we have $F\gg 1$ so that the effective mass of $\chi$ is positive and $\chi$ acts as a damped harmonic oscillator (See Fig. 1). In our model, we find that the coupling parameter $\xi$ must satisfy $\xi>3/16$ to ensure that $\chi$ undergoes an under-damped oscillators during St-1 so that it settles down to the stable local minimum $\chi_{0}$ in a few e-folds. This secures that St-1 can be effectively treated as single-field slow-roll inflation.
@@ 4.3 Superhorizon evolution and $\delta N$
$\mathcal{P}_{\mathcal{R}}(k)\approx\frac{M^{2}}{16\pi^{2}M_{\rm pl}^{2}}\times\left\{\begin{aligned} &\quad\left[\frac{2}{3}\left(\ln\frac{k_{1}}{k}+\frac{3}{4}F_{\star 1}\right)^{2}+g_{1}^{2}h^{2}\left(\frac{M_{\rm pl}}{\chi_{0}}\right)^{2}\frac{H_{\rm exit}}{H_{\star 2}}\left(\frac{k}{k_{1}}\right)^{3}\right]\\ &\qquad\times\left[1-\left(\frac{4}{3}\ln\frac{k_{1}}{k}+F_{\star 1}\right)^{-1}\right]^{2}\quad\mathrm{for}\ k<k_{1}\,,\\ &\quad\frac{g_{2}^{2}h^{2}}{\mu^{2}}\left(\frac{M_{\rm pl}}{\chi_{0}}\right)^{2}\left(\frac{k}{k_{2}}\right)^{-\beta}\quad\mathrm{for}\ k>k_{1}\,,\end{aligned}\right.$ (4.31)
where we have used $H^{2}=M^{2}(1-F^{-1})^{2}/4$ during St-1, $H^{2}=V_{0}/(3M_{\mathrm{pl}})^{2}$ during St-3, $F_{\star 2}=1$, and introduced
@@ 5.1 Primordial black hole formation
Figure 12: Three numerical result of power spectrum $\mathcal{P}_{\mathcal{R}}$ and $f_{\mathrm{PBH}}$. Left panel: We set $\mathcal{P}_{\mathcal{R}}=2.3\times 10^{-9}$ at pivot scale $k=0.002\mathrm{Mpc^{-1}}$ according to [1]. From right to left, the green, blue, and pink curves are numerically calculated with parameter cases 1,2,3 appearing in Table. 1. We tune the parameters and obtain case 1 for PBH as a dominant candidate of cold dark matter, case 2 for the production of sub-solar PBH, and case 3 for explanation of the recently released PTAs observation intentionally. Right panel: From right to left, the green, blue, and pink curves are numerically calculated with parameter case 1,2,3 in Table.1. The corresponding power spectrum is shown by the same color in the left panel. The $f_{\rm PBH}$ constraint plot and codes are available at PBH bounds, we include the observational constraints from PBH evaporation [160] (evap), HSC [161]/EROS [162]/OGLE [17]/MACHO [163]/Kepler [164] microlensing (HSC/EROS/OGLE/M/K) observations and millilensing observation of compact radio sources [165] (Radio), dynamical heating of ultra-faint dwarf galaxies [166] (UFdwarfs), the accretion constraints by CMB anisotropies measured by Planck [167] (CMB), and background constraints by LIGO [168, 169, 170] (LIGO/LIGO-SGWB). We added the latest released 20-year OGLE bounds (OGLE Strict) derived in [171, 172].
@@
$\chi_{0}/M$ $0.275\sqrt{1/(2\pi^{2})}$ $0.172\sqrt{1/(2\pi^{2})}$ $0.145\sqrt{1/(2\pi^{2})}$
@@ 6 Conclusion and Discussion
In terms of the $\delta N$ formalism, this enhancement may be understood as a consequence of the sharp turn of the inflationary trajectory in the field space, an almost orthogonal turn from the $\phi$-direction to the $\chi$-direction. This converts the original isocurvature perturbation due to $\delta\chi$ at St-1 to an adiabatic perturbation at St-3. Since $\delta\chi$ decays like $\sim a^{-3/2}$ during St-1, the isocurvature effect is negligible at early stages. Conversely, it means that the isocurvature effect grows like $a^{3/2}$ toward the end of St-1. This leads to the growth of the curvature perturbation spectrum proportional to $k^{3}$. This is our new finding. This growth rate is universal as long as the isocurvature field ($\chi$ in our case) is heavy enough at the first inflationary stage and dominates the second inflationary stage as a slow-rolling inflaton.
@@
In this paper, we assumed that the curvature perturbation is Gaussian. However, it is known that non-Gaussianities may have a significant influence on the PBH abundance [181, 182, 183, 184, 185, 186] as well as on the PBH clustering [187, 188, 189], which may also leave characteristic features on the corresponding induced GWs [190, 71, 191, 192, 193, 194, 195, 190]. In our model, we found that the non-Gaussianities are of local type, and those generated around the peak of the power spectrum are very small, as expected in the case of multi-field inflationary scenarios with non-extreme transitions [173, 174].